# Vibe AIGC: A New Paradigm for Content Generation via Agentic Orchestration

### v1

@@ 5.4 Agentic Orchestration
Upon the completion of intent expansion and disambiguation, the system enters the stage of Agentic Orchestration. The role of the Meta Planner shifts from creative director to dynamic compiler. It constructs the system by mapping the aforementioned creative scripts into specific algorithmic execution paths, based on the algorithmic workflows and tool definitions stored in the knowledge base Li et al. (2024).
@@ 6 Limitations
The “Bitter Lesson” and Model Centrality. The “Intent-Execution Gap” is not a permanent architectural flaw, but a temporary symptom of insufficient model scale Sutton (2019). If a single foundational model eventually achieves a near-perfect internal world model, the need for a complex orchestration layer may vanish. In this view, “Vibe” is merely a high-entropy prompt that current models cannot yet parse, but future models will execute in a single shot without the overhead of multi-agent delegation.
The Paradox of Control: Commander vs. Craftsman. The shift from prompt engineer to Commander assumes that users prefer high-level intent over granular manipulation. However, professional creators often require “pixel-perfect” control that natural language may inherently lack. Critics argue that Vibe AIGC risks introducing a “Black Box of Intent”. By abstracting away the “How”, we may be trading professional precision for amateur convenience, potentially leading to a “homogenization of aesthetic” where the AI’s interpretation of a vibe overrides the human’s unique creative signature Flusser (2013); Benjamin (2018).
The Verification Crisis: Binary Success vs. Aesthetic Subjectivity. A fundamental challenge for Vibe AIGC is the absence of a deterministic feedback loop. In coding, code either compiles and passes unit tests, or it fails; this verification allows LLMs to iteratively converge on a ground-truth solution Xu et al. (2025). In contrast, a “Vibe” in AIGC is inherently subjective and lacks a formal specification. There is no universal unit test for a “cinematic atmosphere” or “melancholic pacing”. Without an objective verification oracle, the recursive orchestration layer may drift into “aesthetic hallucination” and fail to meet the user’s unstated creative intent.
Compounding Failures and the Missing “Compiler”. The reliance on “recursive orchestration” introduces significant systemic risks related to error compounding. In Coding, a compiler acts as a hard constraint that intercepts logical errors. However, Vibe AIGC relies on multiple agents where a minor semantic drift in an upstream agent can lead to catastrophic hallucinations across the entire workflow Cemri et al. (2025). Unlike modular software, generative artifacts often suffer from “content leakage” or pixel misalignment that current orchestration layers cannot formally “debug”. Critics maintain that until an equivalent of an “Aesthetic Compiler” is developed, multi-agent workflows will remain a fragile compass for digital construction.

### v2

@@ 5.4 Agentic Orchestration
Upon the completion of intent expansion and disambiguation, the system enters the stage of Agentic Orchestration. The role of the Meta Planner shifts from creative director to dynamic compiler. It constructs the system by mapping the aforementioned creative scripts into specific algorithmic execution paths, based on the algorithmic workflows and tool definitions stored in the knowledge base (Li et al., 2024).
@@ 6 Limitations
The “Bitter Lesson” and Model Centrality. The “Intent-Execution Gap” is not a permanent architectural flaw, but a temporary symptom of insufficient model scale (Sutton, 2019). If a single foundational model eventually achieves a near-perfect internal world model, the need for a complex orchestration layer may vanish. In this view, “Vibe” is merely a high-entropy prompt that current models cannot yet parse, but future models will execute in a single shot without the overhead of multi-agent delegation.
The Paradox of Control: Commander vs. Craftsman. The shift from prompt engineer to Commander assumes that users prefer high-level intent over granular manipulation. However, professional creators often require “pixel-perfect” control that natural language may inherently lack. Critics argue that Vibe AIGC risks introducing a “Black Box of Intent”. By abstracting away the “How”, we may be trading professional precision for amateur convenience, potentially leading to a “homogenization of aesthetic” where the AI’s interpretation of a vibe overrides the human’s unique creative signature (Flusser, 2013; Benjamin, 2018).
The Verification Crisis: Binary Success vs. Aesthetic Subjectivity. A fundamental challenge for Vibe AIGC is the absence of a deterministic feedback loop. In coding, code either compiles and passes unit tests, or it fails; this verification allows LLMs to iteratively converge on a ground-truth solution (Xu et al., 2025). In contrast, a “Vibe” in AIGC is inherently subjective and lacks a formal specification. There is no universal unit test for a “cinematic atmosphere” or “melancholic pacing”. Without an objective verification oracle, the recursive orchestration layer may drift into “aesthetic hallucination” and fail to meet the user’s unstated creative intent.
Compounding Failures and the Missing “Compiler”. The reliance on “recursive orchestration” introduces significant systemic risks related to error compounding. In Coding, a compiler acts as a hard constraint that intercepts logical errors. However, Vibe AIGC relies on multiple agents where a minor semantic drift in an upstream agent can lead to catastrophic hallucinations across the entire workflow (Cemri et al., 2025). Unlike modular software, generative artifacts often suffer from “content leakage” or pixel misalignment that current orchestration layers cannot formally “debug”. Critics maintain that until an equivalent of an “Aesthetic Compiler” is developed, multi-agent workflows will remain a fragile compass for digital construction.